# PADetBench: Towards Benchmarking Physical Attacks against Object Detection

### v1

@@ Abstract
Physical attacks against object detection have gained increasing attention due to their significant practical implications. However, conducting physical experiments is extremely time-consuming and labor-intensive. Moreover, physical dynamics and cross-domain transformation are challenging to strictly regulate in the real world, leading to unaligned evaluation and comparison, severely hindering the development of physically robust models. To accommodate these challenges, we explore utilizing realistic simulation to thoroughly and rigorously benchmark physical attacks with fairness under controlled physical dynamics and cross-domain transformation. This resolves the problem of capturing identical adversarial images that cannot be achieved in the real world. Our benchmark includes 20 physical attack methods, 48 object detectors, comprehensive physical dynamics, and evaluation metrics. We also provide end-to-end pipelines for dataset generation, detection, evaluation, and further analysis. In addition, we perform 8064 groups of evaluation based on our benchmark, which includes both overall evaluation and further detailed ablation studies for controlled physical dynamics. Through these experiments, we provide in-depth analyses of physical attack performance and physical adversarial robustness, draw valuable observations, and discuss potential directions for future research. Codebase: GitHub Repository.
@@ 1 Introduction
Deep neural networks (DNNs) have achieved remarkable success in various research fields and industrial applications, such as computer vision [65], natural language processing [64], and speech recognition [61]. However, previous studies [78, 27, 2, 41, 3] have shown that DNNs are susceptible and vulnerable to adversarial attacks, which can be divided into digital attacks and physical attacks according to different attack domains in computer vision tasks. Technically, digital attacks are usually performed by adding imperceptible perturbations to the input image of victim models after imaging, while physical attacks can cheat DNNs by directly modifying the physical properties of the targets before imaging, such as changing the texture of the object [77, 114], and adding stickers [96, 44]. Intuitively, physical attacks are more dangerous and practical than digital attacks because they can be easily implemented in the real world with fewer prerequisites, which arouses significant concerns, especially in safety-critical applications like autonomous driving [92, 6], security surveillance [63, 95], remote sensing [94, 48], and so on.
Object detection is a fundamental and pragmatic task in computer vision, which has been widely deployed in various intelligent systems [126, 113]. Consequently, most of the physical attacks aim to fool object detectors in real-world scenarios, and the physical adversarial robustness of object detection models has attracted increasing attention in recent years. However, the absence of regulated and easy-to-follow benchmarks hinders the development of physical attack and physically robust detection methods. The main reasons for discouraging the development of physical attack benchmarks are as follows. 1) Time-consuming and expensive: To evaluate the physical attack performance and physical adversarial robustness of object detection models, researchers need to conduct a large number of experiments in the real world, which is time-consuming and expensive. 2) Physical dynamics align: To ensure comparison fairness, physical dynamics should be strictly controlled and kept the same, while this can’t be achieved in real-world scenarios since it is impossible to take two identical pictures. 3) Cross-domain loss: Physical attacks are commonly conducted by elaborating conspicuous adversarial patterns that should be able to survive after transforming from the physical to the digital domain, while this cross-domain loss is uncontrollable. 4) Hard to compare: With the progressive evolution of physical attacks, physical adversarial perturbations have evolved from 2D space to 3D space, which builds a barrier to fairly compare different kinds of physical attack methods. As the above challenges exist, it is tough to effectively verify the physical attack efficacy and physical adversarial robustness of object detection models without thorough evaluation and impartial comparisons. As a result, the researchers can not identify the actual progress of physical adversarial attacks and physical adversarial robustness development, which may gear down the progress of the research field.
In this paper, we propose a fair and rigorous benchmarking of physical attacks on object detection models. We utilize CARLA [17], a highly realistic autonomous driving simulator, to benchmark physical attacks under consistent physical dynamics such as weather, viewing angle, and distance. These conditions are challenging to align for impartial comparisons in the real world due to the impossibility of capturing two identical photographs. Our benchmark includes 20 physical attack methods, clean and random noised objects for comparison, 48 object detectors, diverse physical dynamics, evaluation metrics from different perspectives, and a comprehensive pipeline for evaluating physical attacks and subsequent analysis. We conduct exhaustive evaluation of each physical attack against a wide range of object detectors under various physical dynamic conditions, using multiple evaluation metrics. Following our experimental procedures, we delve into detailed analyses concerning the performance of physical attacks and the robustness against physical adversarial actions. We highlight intriguing findings and discuss potential avenues for future research.
In summary, our contributions are as follows:
@@
We propose a robust and equitable benchmark for physical attacks against object detection models. This benchmark deeply explores the potential of real-world simulators to consistently evaluate physical attacks under a variety of continuous physical dynamics.
We offer 8064 groups of evaluation on 20 physical attacks against 48 object detectors, which are conducted under continuous physical dynamics with diverse evaluation metrics, providing a holistic perspective on the performance of attack and detection.
We delve into detailed analyses of the fooling efficacy of physical attacks and the robustness of object detection models against these attacks. Through these analyses, we highlight valuable findings and discuss potential avenues for future research in this area.
@@ 2.1 Object detection
Object detection serves as a cornerstone in computer vision, with the primary objective of identifying and localizing objects within images or video sequences. With the advent of deep learning, DNNs have been widely adopted for object detection, significantly improving performance. R-CNN [26] and its successors [25, 70, 56, 66, 99, 106, 75], have made substantial strides in improving detection speed and accuracy by introducing region proposal networks and shared convolution computations. Single Shot MultiBox Detector (SSD) [55] and You Only Look Once (YOLO) [67] series [67, 68, 69, 1, 38, 43, 86, 37, 87, 85] further accelerate the detection process by eliminating the need for region proposals and performing detection in a single pass, making them suitable for real-time applications. Recently, more advanced architectures based on transformers have been proposed, such as DETR [7], DAB-DETR [52], ViTDet [47], DINO [104], and Co-DETR [125], which further push the boundaries of object detection performance by leveraging attention mechanisms. Despite these advancements, object detection in adversarial environments remains a challenging task, necessitating continuous research and development in this area.
@@ 2.2 Physical attack
Kurakin et al. [41] first illustrate that machine learning systems remain susceptible to adversarial examples, even within these physical world contexts. They substantiate this by introducing adversarial images, captured via a cell phone camera, which still can dramatically degrade the performance of intelligent vision systems. Brown et al. [2] first design adversarial perturbations in patch form, i.e., adversarial patch, which localizes adversarial perturbations to a specific region of the image without imperceptibility constraints. Adversarial patches are more practical and effective in the real world, as they can be easily printed and attached to objects to fool detectors [73, 81, 100, 124, 122, 91, 121, 34, 112, 72, 35, 29]. Nonetheless, the adversarial patches mentioned above are overly noticeable, potentially arousing the suspicion of the victims. To mitigate this concern, scholars have suggested the generation of adversarial patches in natural styles to circumvent drawing attention [36, 32, 28]. Beyond adversarial patches, the academic community has also investigated physical perturbations in a multitude of forms, including light [31, 98], viewpoint [15], and 3D objects [51], among others. To reconcile the disparity between 2D perturbations and 3D physical space, the concept of adversarial perturbations has been expanded into the 3D realm [111, 88, 76, 77, 117]. This extension has proven to be more efficacious and applicable in real-world scenarios. The considerable variation in physical perturbations and experimental settings presents a hurdle to the equitable comparison of disparate physical attack methodologies.
@@ 2.3 Robustness benchmark
The benchmarking of adversarial attacks is crucial for evaluating the robustness of DNNs-based models and further developing more robust ones. To address the issue of overestimated robustness, [12] first establish a standardized benchmark of adversarial robustness, which accurately reflects the robustness of the considered models within a reasonable computational budget. Wu et al. [97] build a comprehensive benchmark for backdoor attacks in image classification models. [60] provides an easy-to-use benchmark to assess the performance of object detection models under conditions of deteriorating image quality, such as image distortions or diverse weather conditions instead of adversarial perturbations. In contrast, [115] benchmark the adversarial robustness of image classifiers in black-box settings. Except for RGB images, Dong et al. [14] benchmark the robustness of 3D object detection to common corruptions from weather, sensor, motion, etc., in both LiDAR and camera data. Li et al. [45] towards benchmarking and assessing the visual naturalness of physical adversarial perturbations. Hingun et al. [30] construct a large-scale benchmark for evaluating adversarial patches with traffic sign dataset. CARLA [17], as a realistic autonomous driving simulator, has been progressively involved in the research of physical adversarial robustness. [62] presents CARLA-GEAR: a dataset generator for evaluating the adversarial robustness of vision models. The work in [109] put forth a pipeline for instance-level data generation, leveraging the CARLA simulator. Utilizing this pipeline, they created the DCI dataset and carried out experiments involving three detectors and three physical attacks. In summary, there is still a lack of a comprehensive and rigorous benchmark for physical attacks against object detection models. We aim to fill the gap in this work and significantly lower the threshold of this path with easy-to-follow instructions and codebase.
@@ 3 PADetBench
The benchmark encompasses four integral facets: datasets, physical attack implementations, object detection models, and comprehensive evaluation & analysis procedures. From a technical standpoint, we have engineered each constituent of the benchmark as modular, end-to-end pipelines within our publicly accessible codebase, ensuring straightforward adoption and replication.
3.1 Dataset Generation
@@
It is common to use COCO [50], PASCAL VOC [19], KITTI [23], etc., as benchmark datasets for object detection. However, these datasets are ill-suited for assessing physical attacks since they are static and lack the flexibility required to create manipulated, real-world adversarial scenarios. Physical attacks typically entail altering the physical attributes of objects before capturing their images. To fairly and accurately evaluate and compare such attacks, experiments necessitate applying perturbations in real-world conditions with controlled physical dynamics, which are excessively time-consuming, labor-intensive, and theoretically infeasible. Simulated environments, like CARLA [17], present a viable solution to these obstacles by enabling the straightforward manipulation of physical dynamics through configurable parameters.
@@
This work contributes an end-to-end pipeline for dataset generation within our codebase, significantly streamlining the dataset generation process and enhancing research productivity. Our pipeline prioritizes user-friendliness, enabling researchers to swiftly generate datasets embodying diverse physical conditions through a concise series of steps. These conditions encompass variations in weather, viewing angles, and distances, along with the capacity to impose physical perturbations on objects. Comprehensively, our pipeline supports over 10 distinct environments ranging from downtowns to small towns and rural landscapes (detailed listing available in the Appendix), coupled with a library of more than 40 vehicles and 40 pedestrian models, all customizable concerning their hues and surface textures. It further integrates continuous manipulation of physical dynamics such as fluctuating weather patterns, precise sun positioning, and flexible camera placements concerning both location and orientation. To ensure accessibility, we accompany the pipeline with step-by-step guidelines for personalizing object perturbations and seamlessly integrating these modifications within CARLA’s [17] simulation framework.
@@
Our benchmark comprises three categories of datasets: a clean dataset serving as a control group, a dataset with random noise perturbations, and several datasets featuring adversarial perturbations generated through various attack methodologies. To ensure fair comparisons, scene compositions and camera perspectives are meticulously synchronized and regulated across all datasets, achievable effortlessly through our provided pipeline.
@@
Moreover, our pipeline facilitates the automatic generation of supplementary annotations, including 2D and 3D bounding boxes, depth maps, and instance segmentation maps. Consequently, our benchmark extends its utility beyond 2D object detection, also catering to tasks like 3D object detection, instance segmentation, depth estimation, and more, thereby enhancing the scope of research and application in computer vision.
@@ 3.2 Physical attacks
We select 20 representative physical attack methods, in which FCA [88], DTA [76], ACTIVE [77], 3D2Fool [114], POOPatch [11], RPAU [54], APPA [48], and CAMOU [111], are for vehicle detection, while DAP [29], AdvPattern [95], UPC [36], NatPatch [32], MTD [13], AdvCaT [33], AdvTexture [34], AdvTshirt[101], AdvPatch [81], LAP [79], InvisCloak [100], and AdvCam [18], are for person detection. The corresponding physical perturbations of these methods are imported into Unreal Engine 4 for CARLA (CarlaUE4) [17], as shown in the physical attacks part of Fig. 1, to generate the physical adversarial datasets. We adhere to two principles similar to [97] when selecting physical attacks. First, the methods are representative or advanced in the research field, which can serve as baseline and state-of-the-art (SOTA) methods for comparison, respectively. Second, physical attacks are easily conducted and with reproducible performance, which can be conveniently followed and reproduced by other researchers. Since our benchmark evaluates physical attacks based on their crafted perturbations, novel physical attack methods can be easily integrated into the benchmark by following the provided pipeline. We will continue to update the physical attacks in the benchmark to keep pace with the latest research progress.
@@
CNN | One-stage | ATSS[107], AutoAssign[119], CenterNet[118], CentripetalNet[16] |
CornerNet[42], DDOD[10], DyHead[99], EfficientNet[80], FCOS[82], | ||
FoveaBox[40], FreeAnchor[110], FSAF[120], GFL[46],LD[116], NAS-FPN[24], | ||
PAA[39], RetinaNet[49], RTMDet[58], TOOD[20], VarifocalNet[105], | ||
YOLOX[22], YOLOv5[38], YOLOv6[43], YOLOv7[86], YOLOv8[37] | ||
Two-stage | Faster R-CNN[70], Cascade R-CNN[5], Cascade RPN[84], | |
Double Heads[99], FPG[8],Grid R-CNN[56], Guided Anchoring[89], | ||
HRNet[74], Libra R-CNN[66], PAFPN[53],RepPoints[102], | ||
Res2Net[21], ResNeSt[103], SABL[90], Sparse R-CNN[75] | ||
Transformer | - | DETR[7], Conditional DETR[59], DDQ[108], DAB-DETR[52], |
Deformable DETR[123], DINO[104],PVT[93], PVTv2[93] |

### v3

@@ Abstract
Physical attacks against object detection have gained significant attention due to their practical implications. However, conducting physical experiments is time-consuming and labor-intensive, and controlling physical dynamics and cross-domain transformations in the real world is challenging, leading to inconsistent evaluations and hindering the development of robust models. To address these issues, we rigorously explore realistic simulations to benchmark physical attacks under controlled conditions. This approach ensures fairness and resolves the problem of capturing strictly aligned adversarial images, which is challenging in the real world. Our benchmark includes 23 physical attacks, 48 object detectors, comprehensive physical dynamics, and evaluation metrics. We provide end-to-end pipelines for dataset generation, detection, evaluation, and analysis. The benchmark is flexible and scalable, allowing easy integration of new objects, attacks, models, and vision tasks. Based on this benchmark, we generate comprehensive datasets and perform over 8,000 evaluations, including overall assessments and detailed ablation studies. These experiments provide detailed analyses from detection and attack perspectives, highlight limitations of existing algorithms, and offer revealing insights. The code and datasets are publicly available at https://github.com/JiaweiLian/PADetBench.
keywords:
[…]
@@ 1 Introduction
Deep neural networks (DNNs) have achieved remarkable success in various fields such as computer vision, natural language processing, and speech recognition. However, studies [1, 2, 3, 4] show that DNNs are vulnerable to adversarial attacks, which can be categorized into digital and physical attacks. Digital attacks add imperceptible perturbations to input images post-imaging, while physical attacks modify the physical properties of targets pre-imaging, such as changing textures [5, 6] or adding stickers [7, 8]. Physical attacks are more practically dangerous as they can be easily implemented in real-world scenarios, raising significant concerns in safety-critical applications like autonomous driving [9, 10, 11], security surveillance [12, 13], and remote sensing [14, 15].
Object detection is a fundamental and pragmatic task in computer vision, widely deployed in various intelligent systems [16, 17, 18, 19]. Consequently, many physical attacks aim to fool object detectors in real-world scenarios, and the physical adversarial robustness of object detection models has garnered increasing attention in recent years. However, the absence of regulated and easy-to-follow benchmarks hinders the development of physical attack and physically robust detection methods. The main reasons for the lack of physical attack benchmarks are concluded as follows:
Time-consuming and expensive: Evaluating the performance of physical attacks and the adversarial robustness of object detection models requires numerous real-world experiments, which are both time-consuming and costly.
Physical dynamics alignment: Ensuring fair comparisons necessitates strictly controlled and consistent physical dynamics, which is difficult to achieve in real-world scenarios where it is nearly impossible to capture identical physical dynamics.
Cross-domain loss: Physical attacks often involve creating conspicuous adversarial perturbations that must survive the transformation from the physical to the digital domain and vice versa. This cross-domain loss is often uncontrollable and unpredictable.
Difficulty in comparison: With the evolution of physical attacks from 2D to 3D space, it becomes increasingly challenging to compare different types of physical attack methods fairly.
Due to these challenges, it is difficult to effectively verify the efficacy of physical attacks and the adversarial robustness of object detection models without thorough evaluation and impartial comparisons. As a result, researchers cannot accurately gauge the progress of physical adversarial attacks and robustness development, which slows down advancements in the field.
This paper proposes utilizing realistic simulations to benchmark physical attacks under controlled conditions such as weather, viewing angle, and location. These conditions are challenging to align for impartial comparisons in the real world. Our benchmark includes 23 physical attack methods, 48 object detectors, diverse physical dynamics, evaluation metrics from different perspectives, comprehensive pipelines for data generation, attack and detection evaluation, and subsequent analysis. Moreover, the benchmark is highly flexible and scalable, allowing for easy integration of new physical attacks, models, and other vision tasks. Based on the benchmark, we generate comprehensive and strictly aligned datasets and perform over 8,000 evaluations, including overall assessments and detailed ablation studies for controlled physical dynamics. Through these experiments, we provide detailed analyses from detection and attack perspectives, highlight algorithm limitations, and convey valuable insights. In summary, our contributions are as follows:
@@
We propose a robust and equitable benchmark for physical attacks against object detection models. This benchmark deeply explores the potential of real-world simulators to evaluate physical attacks under various continuous and strictly aligned physical dynamics.
The benchmark includes 23 physical attacks, 48 object detectors, comprehensive physical dynamics, and rigorous evaluation metrics. We provide end-to-end pipelines for dataset generation, detection, evaluation, and analysis, ensuring a thorough evaluation process.
The benchmark is designed to be highly flexible and scalable, facilitating the easy integration of new objects, attacks, detectors, and even other vision tasks. This adaptability enhances the utility of our framework for ongoing research and development in the field.
Based on our benchmark, we generate comprehensive datasets and perform over 8,000 evaluations, including overall assessments and detailed ablation studies. These experiments highlight the limitations of existing algorithms and illuminate informative insights.
@@ 2.1 Object detection
Object detection is a fundamental task in computer vision, aiming to identify and localize objects within images or videos. It can be formulated as a mapping function , where is the input space and is the output space (e.g., bounding boxes and class labels). Deep learning has significantly advanced object detection. R-CNN [20] and its successors [21, 22, 23, 24, 25, 26, 27] improved detection speed and accuracy with region proposal networks and shared convolution computations. SSD [28] and YOLO series [29, 30, 31, 32, 33, 34, 35, 36, 37, 38] further accelerated detection by eliminating region proposals, enabling real-time applications. Recently, transformer-based architectures like DETR [39], DAB-DETR [40], ViTDet [41], DINO [42], and Co-DETR [43] have pushed performance boundaries using attention mechanisms. Despite increasing advancements [44, 45, 46], object detection in adversarial environments remains challenging, requiring ongoing research.
@@ 2.2 Physical attack
Adversarial attacks typically add imperceptible perturbations to the clean input in the digital domain, fooling DNNs into incorrect predictions. This is formulated as: , where is the attack loss and is the ground-truth. In contrast, physical attacks often manipulate the physical properties of objects to deceive detection models, formulated as: , where and are transformations between digital and physical domains. [47] first showed that machine learning systems are vulnerable to adversarial examples in physical contexts. They demonstrated this with adversarial images captured via a cell phone camera, significantly degrading vision system performance. [48] introduced adversarial patches, which localize perturbations to specific image regions without imperceptibility constraints. These patches are practical and effective in the real world, easily printed and attached to objects to fool detectors [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60]. To avoid suspicion, natural-style adversarial patches have been proposed [61, 62, 63]. Beyond patches, physical perturbations include light [64, 65], viewpoint [66], and 3D objects [67]. Extending adversarial perturbations to 3D space [68, 69, 70, 5, 71] has proven more effective and applicable in real-world scenarios. The variety in perturbations and settings complicates fair comparisons of physical attack methods.
@@ 2.3 Robustness benchmark
Benchmarking adversarial attacks is crucial for evaluating and improving the robustness of DNN-based models. [72] established a standardized benchmark for adversarial robustness, accurately reflecting model robustness within a reasonable computational budget. [73] created a comprehensive benchmark for backdoor attacks in image classification models. [74] provided a benchmark to assess object detection models under deteriorating image quality, such as distortions or adverse weather conditions. [75] benchmarked the adversarial robustness of image classifiers in black-box settings. [76] evaluated the robustness of 3D object detection to common corruptions in LiDAR and camera data. [77] focused on benchmarking the visual naturalness of physical adversarial perturbations. [78] constructed a large-scale benchmark for evaluating adversarial patches using a traffic sign dataset. CARLA [79], a realistic autonomous driving simulator, has been widely used in physical adversarial robustness research. [80] presented CARLA-GEAR, a dataset generator for evaluating the adversarial robustness of vision models. [81] proposed a pipeline for instance-level data generation using CARLA, creating the DCI dataset and conducting experiments with three detectors and three physical attacks. Despite these efforts, a comprehensive and rigorous benchmark for physical attacks against object detection models is still lacking. This work aims to fill that gap by providing easy-to-follow instructions and a robust codebase.
@@ 3 PADetBench
The benchmark encompasses four integral facets: dataset generation, physical attacks, object detection, and comprehensive evaluation & analysis procedures, as shown in Fig. 1. From a technical standpoint, we have engineered each benchmark constituent as modular, end-to-end pipelines within the codebase, ensuring straightforward adoption and replication.
3.1 Datasets Generation
@@
It is common to use COCO [82], PASCAL VOC [83], KITTI [84], etc., as benchmark datasets for object detection. However, these datasets are ill-suited for assessing physical attacks since they are static and lack the flexibility to create manipulated, real-world adversarial scenarios. Physical attacks typically entail altering the physical attributes of objects before capturing their images. To fairly and accurately evaluate and compare such attacks, experiments necessitate applying perturbations in real-world conditions with controlled physical dynamics, which are excessively time-consuming, labor-intensive, and theoretically infeasible. Simulated environments, like CARLA [79], offer a viable solution to these obstacles by enabling the straightforward manipulation of physical dynamics through configurable parameters.
@@
This work contributes an end-to-end pipeline for dataset generation within our codebase, significantly streamlining the dataset generation process and enhancing research productivity. Our pipeline prioritizes user-friendliness, enabling researchers to generate datasets embodying diverse physical conditions through concise steps swiftly. These conditions encompass variations in weather, viewing angles, distances, and the capacity to impose physical perturbations on objects. Comprehensively, our pipeline supports over ten distinct environments ranging from downtowns to small towns and rural landscapes. It is coupled with a library of more than 40 vehicles and 40 pedestrian models, all customizable concerning their hues and surface textures. It further integrates continuous manipulation of physical dynamics such as fluctuating weather patterns, precise sun positioning, and flexible camera placements concerning location and orientation.
We provide a detailed illustration of the physical dynamics alignment in Fig. 2 and Fig. 3. Specifically, it is observed from Fig. 2 that the imaging settings and lighting conditions are not strictly aligned in the comparison experiments, such as the different view angles and shadows, which have been demonstrated to have a significant impact on fooling deep neural networks [85, 66]. To address this issue, we align the physical dynamics in the benchmark, as shown in Fig. 3, where the physical dynamics are strictly controlled and aligned, ensuring a fair and impartial comparison. Moreover, we also provide a detailed illustration of the physical dynamics in Fig. 4, which includes the weather conditions, camera settings, and lighting conditions. The lighting conditions vary aligned with real-world laws as shown in Fig. 5, such as the sun positions of 24 hours, the intensity of the light, and the shadow, which are strictly controlled and aligned in the benchmark. In addition, we detail the diverse scenes and objects in A.2.
@@
Our benchmark comprises three datasets: a clean dataset serving as a control group, a dataset with random noise perturbations, and several datasets featuring adversarial perturbations generated through various attack methodologies. To ensure fair comparisons, scene compositions and camera perspectives are meticulously synchronized and regulated across all datasets, achievable effortlessly through our provided pipeline. To ensure accessibility, we accompany the pipeline with step-by-step guidelines for personalizing object perturbations and seamlessly integrating these modifications within CARLA’s [79] simulation framework.
@@
Moreover, our pipeline facilitates the automatic generation of supplementary annotations, including 2D and 3D bounding boxes, depth maps, and instance segmentation maps. Consequently, our benchmark extends its utility beyond 2D object detection, catering to tasks like 3D object detection, instance segmentation, depth estimation, and more, thereby enhancing the scope of research and application in computer vision.
@@ 3.2 Physical attacks
Physical attacks are usually tailored for specific objects, and the commonly targeted objects are vehicles, persons, and traffic signs, as evidenced by [86]. Consequently, we adopt typical objects from these categories as examples to illustrate the proposed benchmark. Specifically, we select 23 representative physical attack methods, categorized into three types according to their target objects: vehicle, person, and traffic sign, as shown in Table 1. The corresponding physical perturbations of these methods are imported into Unreal Engine 4 for CARLA (CarlaUE4) [79], as shown in the physical attacks part of Fig. 1, to generate the physical adversarial datasets. When selecting physical attacks, we adhere to two principles similar to [73]. First, the methods are representative or advanced in the research field, which can serve as a baseline and state-of-the-art (SOTA) methods for comparison, respectively. Second, physical attacks are efficiently conducted with reproducible performance, which other researchers can conveniently follow and reproduce. Since our benchmark evaluates physical attacks based on their crafted perturbations, novel physical attack methods can be easily integrated into the benchmark by following the provided pipeline. We will continue to update the physical attacks in the benchmark to keep pace with the latest research progress.
Target objects | Physical attacks |
Vehilcle | FCA [69], DTA [70], ACTIVE [5], |
3D2Fool [6], POOPatch [87], | |
RPAU [88], CAMOU [68] | |
Person | DAP [60], AdvPattern [13], UPC [61], |
NatPatch [62], MTD [89], AdvCaT [90], | |
AdvTexture [56], AdvTshirt[91], AdvPatch [50], | |
LAP [92], InvisCloak [51], AdvCam [93] | |
Traffic sign | AdvCam [93], RP2 [94], ShapeShifter[95] |
@@
CNN | One-stage | ATSS[96], AutoAssign[97], NAS-FPN [98] |
LD[99], GFL[100], CornerNet[101], PAA[102], | ||
DDOD[103], DyHead[25], EfficientNet[104], | ||
FCOS[105], FoveaBox[106], FreeAnchor[107], | ||
CenterNet[108], CentripetalNet[109], FSAF[110], | ||
RTMDet[111], TOOD[112], VarifocalNet[113], | ||
YOLOX[114], YOLOv5[33], YOLOv6[34], | ||
YOLOv7[35], RetinaNet[115], YOLOv8[36] | ||
Two-stage | Faster R-CNN[22], Cascade R-CNN[116], | |
Cascade RPN[117], Double Heads[25], FPG[118], | ||
Libra R-CNN[24], PAFPN[119], HRNet[120], | ||
ResNeSt[121], Res2Net[122], SABL[123], | ||
Guided Anchoring[124], Sparse R-CNN[27], | ||
RepPoints[125], Grid R-CNN[23] | ||
Transformer | - | DETR[39], PVT[126], PVTv2[126], |
DDQ[127], DAB-DETR[40], DINO[42], | ||
Deformable DETR[128], Conditional DETR[129] |
Clean |
Random |
ACTIVE |
DTA |
FCA |
APPA |
POOPatch |
3D2Fool |
CAMOU |
RPAU |
Please note that the categorization is based on the selected version of the object detection methods, which means the category may vary with their different versions, such as the backbone of a detector can be CNN or Transformer. Please refer to A.3 for the corresponding config files of the detectors.